# Protecting Deep Learning Model Copyrights with Adversarial Example-Free Reuse Detection

### v1

@@ Abstract
Model reuse techniques can reduce the resource requirements for training high-performance deep neural networks (DNNs) by leveraging existing models. However, unauthorized reuse and replication of DNNs can lead to copyright infringement and economic loss to the model owner. This underscores the need to analyze the reuse relation between DNNs and develop copyright protection techniques to safeguard intellectual property rights. Existing DNN copyright protection approaches suffer from several inherent limitations hindering their effectiveness in practical scenarios. For instance, existing white-box testing-based approaches cannot address the common heterogeneous reuse case where the model architecture is changed, and DNN fingerprinting approaches heavily rely on generating adversarial examples with good transferability, which is known to be challenging in the black-box setting. To bridge the gap, we propose NFARD, a Neuron Functionality Analysis-based Reuse Detector, which only requires normal test samples to detect reuse relations by measuring the models’ differences on a newly proposed model characterization, i.e., neuron functionality (NF). A set of NF-based distance metrics is designed to make NFARD applicable to both white-box and black-box settings. Moreover, we devise a linear transformation method to handle heterogeneous reuse cases by constructing the optimal projection matrix for dimension consistency, significantly extending the application scope of NFARD. To the best of our knowledge, this is the first adversarial example-free method that exploits neuron functionality for DNN copyright protection. As a side contribution, we constructed a reuse detection benchmark named Reuse Zoo that covers various practical reuse techniques and popular datasets. Extensive evaluations on this comprehensive benchmark show that NFARD achieves F1 scores of 0.984 and 1.0 for detecting reuse relationships in black-box and white-box settings, respectively, while generating test suites $2\sim 99$ times faster than previous methods.
@@ I Introduction
Deep learning has been extremely successful in a wide range of tasks [1, 2, 3]. However, training a high-performance deep neural network (DNN) requires not only a large amount of labeled data but also significant computational resources that escalate with the complexity of the model. For example, it cost more than \$100 million to train GPT-4 [4], one of the most advanced DNNs in existance. Meanwhile, just like traditional software, deep learning models are often reused and modified for another learning task to save resources. To this end, many model reuse techniques have been proposed to avoid training models from scratch. For example, transfer learning [5] reduces the required data by applying knowledge gained on one problem to another different but similar task. Model compression [6] reduces the model size while preserving reasonable performance so that DNNs can be deployed in resource-constrained scenarios.
@@
Despite the benefits, model reuse techniques also pose concerns about copyright infringement. DNN models can be exposed in multiple ways without adequate protection, including distributed copies and Machine Learning as a Service (MLaaS) [7]. Unauthorized reuse, such as commercialization in breach of license or theft via model extraction techniques [8, 9, 10] (a form of model reuse), can lead to copyright infringement, posing threats to the competitive edge of the DL model owner and causing significant economic loss. With the push for AI governance, such as the OECD recommendation on AI [11] and the European Union’s Artificial Intelligence Act [12], DNN copyright protection is becoming increasingly important to the responsible development of deep learning models in both academic and industrial practices. There is a growing need to develop DNN copyright protection methods to prevent the valuable intellectual properties from being illegally replicated and to provide the necessary evidence to resolve copyright infringement disputes.
@@
Several types of DNN copyright protection methods have been proposed, including watermarking, fingerprinting, and testing-based similarity comparison. Watermarking methods [13, 14, 15, 16, 17] embed a secret watermark into the model during training by utilizing over-parameterization. Then the owner can identify whether a suspect model is a copy by comparing the extracted watermarks. However, the watermark needs to be embedded during the training phase, which may compromise the performance and security of the model. Moreover, recent works have shown that adaptive attacks [18, 19] can remove watermarks. Fingerprinting methods [20, 21, 22], on the other hand, do not tamper with the training procedure. Instead, it extracts a fingerprint from the model using a set of test samples, which contains characterizing features of the model that remain unchanged after applying DNN reuse techniques. Any model with a similar fingerprint will be judged as a copy. Most existing fingerprinting methods heavily rely on generating adversarial examples with good transferability, which is known to be challenging, especially in the black-box setting [23]. Another line of testing-based similarity comparison approaches [24, 25] employs a set of testing metrics and specially crafted test samples to measure the models’ similarity to detect reuse relations. However, existing white-box testing methods cannot address the common heterogeneous reuse case where the model architecture is changed. In summary, existing DNN copyright protection approaches suffer from inherent limitations that hinder their practical use in diverse scenarios.
@@
To address the gap, we propose NFARD, a Neuron Functionality Analysis-based Reuse Detector, which only requires normal test samples to detect reuse relations by measuring the models’ differences on a newly proposed model characterization, i.e., neuron functionality (NF). Neuron functionality has been well practiced to characterize a DNN in neural network representation studies [26, 27]. In this work, we leverage neuron functionality to protect DNN copyright. Specifically, a neuron’s functionality is the mapping from the input space to the one-dimensional output space of the neuron, i.e., $\mathcal{X}\rightarrow\mathbb{R}$ with $\mathcal{X}$ representing the input space. [^1] Neuron functionality is the direct mapping function which is fundamentally different from neuron output statistics used in previous approaches. Recent studies [26, 27, 28] have shown that independently trained neural networks have divergent neuron functionalities. On the other hand, neural networks obtained by reuse techniques have similar neuron functionalities to those of the original model. Therefore, we can detect model reuse relations by measuring the similarity of neuron functionalities, which can be approximated only using normal samples, avoiding the trouble of carefully generating fingerprints or test samples. We devise a set of distance metrics to measure the similarity of neuron functionality under both white-box and black-box settings. Besides, in order to extend the application scope of NFARD, we further propose a linear transformation method to map the neuron functionality of neural networks with different architectures or classification tasks to a common feature dimension. The key is to identify the optimal projection matrix from high-dimensional features to a low-dimensional model representation. We compute the optimal projection by solving a set of least squares problems. This enables NFARD to effectively address heterogeneous cases, facilitating its utility in diverse practical scenarios.
@@
To enable a comprehensive evaluation of NFARD, we construct a model reuse detection benchmark named Reuse Zoo, which consists of 250 neural network models, covering seven commonly used model reuse techniques, three datasets, and four popular model architectures. We use this benchmark to evaluate NFARD and compare it with three baseline approaches: a representative fingerprinting method, IPGuard [20], and two state-of-the-art testing-based similarity comparison approaches, i.e., ModelDiff [25] and DeepJudge [24]. Extensive experimental results show that NFARD can effectively detect reuse relations in both black-box and white-box settings, achieving an F1-score of 0.984 and 1.0, respectively, and is $2\sim 99$ times faster than previous methods in terms of test suite generation.
In summary, our main contributions are as follows:
@@
    We propose NFARD, a neuron functionality analysis-based reuse detector for DNN copyright protection, which incorporates a new model characterization and a set of distance metrics in both white-box and black-box settings. Compared to existing approaches, NFARD does not rely on generating fingerprints or test samples.
@@
    We propose a linear transformation approach via least squares solving to address the heterogeneous reuse case, significantly extending the application scope of the proposed reuse detector.
@@
    We construct and release a model reuse benchmark with 250 models for evaluating DNN copyright protection methods, covering a variety of popular model reuse types.
@@
    We extensively evaluate NFARD to show that it has a broader application scope and can achieve better detection results than SOTA methods, especially on heterogeneous cases.
@@ II-A Deep Neural Network and Neuron Functionality
We focus on classification tasks in this work. A deep neural network classifier is a function mapping from the input space to a probability distribution, i.e., $f:\mathcal{X}\to[0,1]^{m}$ where $\mathcal{X}$ is the input space and $m$ is the number of classes. The neural network includes $L$ layers $f^{(1)},f^{(2)},\cdots,f^{(L)}$ whose output dimensions are $m_{1},m_{2},\cdots,m_{L}$, respectively. We assume $f^{(L)}$ is a linear classifier and call it the last layer of the model. Given an input $\bm{x}\in\mathcal{X}$, the output of $f$ on $\bm{x}$ is defined by $f=\sigma\circ f^{(L)}\circ\cdots\circ f^{(1)}$, where $\sigma$ is the softmax function that converts any vector to a probability vector.
In a neural network $f$, the functionality of a neuron is the mapping from the input space to the one-dimensional output space of the neuron, i.e., $f_{i}^{(k)}:\mathcal{X}\to\mathbb{R}$, where $f_{i}^{(k)}$ is the $i$-th neuron of the $k$-th layer. Mathematically, the functionality of a neuron $f_{i}^{(k)}$ is a highly non-convex function. In practice, it is approximated by a set of normal samples $\{\bm{x}_{j}\}_{j=1}^{n}$ and characterized as a vector $\bm{\mathrm{nv}}(k,i)$:
@@
Recent work [26, 27] on neural network representation proposed a neuron functionality analysis approach to study DNN learning dynamics and interpretability. They have shown that independently trained DNN models with different weight initialization have different corresponding neuron functionalities, which is in accordance with [28]’s conclusions. That is to say, even if two DNN models $f$ and $g$ have the same structure, training dataset, and training setting, their corresponding neuron functionalities $\bm{\mathrm{nv}}_{f}(k,i)$ and $\bm{\mathrm{nv}}_{g}(k,i)$ can vary greatly as long as they have different weight initialization.
@@ II-B Model Reuse Detection for Copyright Protection
Given a victim model $f_{v}$ and a suspect model $f_{s}$, the task of the reuse detector is to determine whether $f_{s}$ is a positive suspect model, i.e., a surrogate model obtained by applying reuse methods to $f_{v}$, or a negative suspect model trained from scratch or irrelevant to $f_{v}$. The reuse detector may require a set of negative suspect models $\{f_{r,i}\}_{i=1}^{R}$ trained independently of the victim model serving as references to make the decision. Such models are called reference models [^3] In practice, hyperparameter tuning process (e.g., grid search) produces multiple independently trained models that can be used as reference models.. Depending on the access of the reuse detector to the models, the detection methods can be divided into two categories: white-box and black-box methods. In the white-box setting, the detector has access to the internals of the neural networks and the final probability output. While in the black-box setting, the detector only has access to the final probability output. In either case, we assume that a subset of the training data of the victim model is available for the reuse detector (e.g., the model owners themselves are searching for surrogate models).
@@ II-C Model Reuse Methods
We consider seven commonly used DNN reuse methods, including fine-tuning, retraining, pruning, quantization, knowledge distillation, transfer learning, and model extraction. These reuse methods are classified into homogeneous and heterogeneous methods depending on whether the obtained surrogate model has the same structure as the victim model.
Fine-tuning and retraining techniques train the victim model for a few epochs with a relatively small learning rate, and some weight parameters may be frozen [30] to enable a finer control. The difference between them is that retraining also re-initializes the weights of the last layer. Pruning and quantization methods are compression techniques that aim to reduce DNN size by removing trivial weights or quantizing parameters to low-bit values while minimizing the drop in model performance [31, 32]. Surrogate models obtained by these four methods have the same model structure as the victim model, so they belong to homogeneous methods. On the other hand, knowledge distillation and model extraction train surrogate models by exploiting the victim model’s outputs [33, 9, 10], and they usually result in a different architecture; transfer learning replaces the last layer with a new one [5], and the classification task becomes different. Therefore, these three reuse techniques change the model structure and fall into the category of heterogeneous methods.
Detecting heterogeneous cases is usually more challenging. To the best of our knowledge, in previous work, only ModelDiff [25] could handle transfer learning reuse cases, but it is not good at dealing with the other two heterogeneous reuse types. We propose a benchmark Reuse Zoo covering different types of reuse methods for DNN reuse detector evaluation, and we elaborate on the construction details in Section IV.
@@ III-A Insight into Reused DNN Models
To begin with, we introduce two metrics to measure the similarity of neuron functionality. The first metric is $dist_{eu}^{(k)}$, which is the average Euclidean distance between the $k$-th layer’s neuron vectors of two models, defined as:
$dist_{eu}^{(k)}(f_{v},f_{s},X)=\dfrac{1}{m_{k}}\sum\limits_{i=1}^{m_{k}}\lVert\bm{\mathrm{nv}}_{v}(k,i)-\bm{\mathrm{nv}}_{s}(k,i)\rVert_{2},$
where $X=\{\bm{x}_{1},\cdots,\bm{x}_{n}\}$ is a set of test inputs, $\bm{\mathrm{nv}}_{v}(k,i)$ is the $i$-th neuron vector of the $k$-th layer of $f_{v}$, and $\bm{\mathrm{nv}}_{s}(k,i)$ is that of $f_{s}$. The second metric is the average cosine distance of the neuron vectors, denoted as $dist_{ac}^{(k)}$:
$dist_{ac}^{(k)}(f_{v},f_{s},X)$ $=\dfrac{1}{m}\sum\limits_{i=1}^{m}dist_{cos}(\bm{\mathrm{nv}}_{v}(k,i),\bm{\mathrm{nv}}_{s}(k,i)),$
where $dist_{cos}$ represents cosine distance. Compared with $dist_{eu}^{(k)}$, $dist_{ac}^{(k)}$ focuses more on the difference in direction rather than the scale. The smaller the metrics, the more similar the two models are in terms of neuron functionality. We do not include metrics used in previous DNN representation studies (such as SVCCA similarity [26]) because they target analyzing training dynamics on hidden layers where correlation analysis is essential, and our goal is to detect reuse relations using the output layer neuron vectors.
@@
Next, we examine four victim models with different structures and learning tasks. For each victim model, we prepare 8 surrogate models by applying homogeneous reuse methods and 40 reference models that are independently trained on the same task (details of these models are available in Section IV). We compute the distance metrics $dist_{eu}^{(L)}$ and $dist_{ac}^{(L)}$ on the last layer (before the softmax function) using 1000 normal samples, and the results are presented in Figure 1, where surrogate models and reference models are plotted in orange and blue, respectively. We can see that reference models have larger distance values and are concentrated in the upper right corner of the plane, while surrogate models have smaller distance values and are scattered on the lower left part. Such differences indicate that homogeneous surrogate models are more similar to victim models than independently trained models regarding neuron functionality, which is exactly the principle of our proposed method.

### v2

@@ Abstract
Model reuse techniques can reduce the resource requirements for training high-performance deep neural networks (DNNs) by leveraging existing models. However, unauthorized reuse and replication of DNNs can lead to copyright infringement and economic loss to the model owner. This underscores the need to analyze the reuse relation between DNNs and develop copyright protection techniques to safeguard intellectual property rights. Existing DNN copyright protection approaches suffer from several inherent limitations hindering their effectiveness in practical scenarios. For instance, existing white-box fingerprinting approaches cannot address the common heterogeneous reuse case where the model architecture is changed, and DNN fingerprinting approaches heavily rely on generating adversarial examples with good transferability, which is known to be challenging in the black-box setting. To bridge the gap, we propose NFARD, a Neuron Functionality Analysis-based Reuse Detector, which only requires normal test samples to detect reuse relations by measuring the models’ differences on a newly proposed model characterization, i.e., neuron functionality (NF). A set of NF-based distance metrics is designed to make NFARD applicable to both white-box and black-box settings. Moreover, we devise a linear transformation method to handle heterogeneous reuse cases by constructing the optimal projection matrix for dimension consistency, significantly extending the application scope of NFARD. To the best of our knowledge, this is the first adversarial example-free method that exploits neuron functionality for DNN copyright protection. As a side contribution, we constructed a reuse detection benchmark named Reuse Zoo that covers various practical reuse techniques and popular datasets. Extensive evaluations on this comprehensive benchmark show that NFARD achieves F1 scores of 0.984 and 1.0 for detecting reuse relationships in black-box and white-box settings, respectively, while generating test suites $2\sim 99$ times faster than previous methods.
@@ I Introduction
Deep learning (DL) has achieved remarkable success across a wide range of tasks [1, 2, 3]. However, training high-performance deep neural networks (DNNs) demands vast amount of labeled data and substantial computational resources, and these costs escalate with model complexity. Training state-of-the-art models like GPT-4 [4], for instance, reportedly cost over \$100 million. Consequently, reusing and adapting existing DL models for new tasks has become common practice to save resources. Techniques such as transfer learning [5], which leverages knowledge from one task for another related one, and model compression [6], which reduces model size for deployment in resource-constrained environments, exemplify this trend towards efficient model reuse.
@@
Despite the benefits, model reuse techniques also pose concerns regarding copyright infringement. DNNs can be exposed in multiple ways without adequate protection, including distributed copies and Machine Learning as a Service (MLaaS) [7]. Unauthorized reuse, such as commercialization in breach of license or theft via model extraction techniques [8, 9, 10] (a form of model reuse), constitutes copyright infringement, posing threats to the competitive edge of the DL model owner and causing significant economic loss. With the push for AI governance, such as the OECD recommendation on AI [11] and the European Union’s Artificial Intelligence Act [12], DNN copyright protection is becoming increasingly important to the responsible development of DL models in both academic and industrial practices. There is a growing need to develop DNN copyright protection methods to prevent the valuable intellectual properties from being illegally replicated and to provide the necessary evidence to resolve copyright infringement disputes.
@@
A number of DNN copyright protection methods have been proposed, including two major types: watermarking and fingerprinting. Watermarking methods [13, 14, 15, 16, 17] embed hidden identifiers into the model’s parameters during training, leveraging over-parameterization. Then the owner can identify whether a suspect model is a copy by comparing the extracted watermarks. However, the watermark needs to be embedded during the training phase, which may compromise the performance and security of the model. Moreover, recent studies have shown that watermarks can be vulnerable to removal via adaptive attacks [18, 19]. In contrast, fingerprinting methods [20, 21, 22, 23, 24] do not tamper with the training procedure. They extract fingerprints from the model using a set of carefully crafted test samples, which contains characterizing features of the model that remain unchanged after applying DNN reuse techniques. Any model with a similar fingerprint is deemed a derivative. However, fingerprinting methods face two main challenges. First, most existing fingerprinting methods depend significantly on the generation of adversarial examples with high transferability, a task that is particularly difficult in black-box settings [25]. Therefore, they mostly assume the availability of gradient information, which may not hold for deployed DL models. Second, current fingerprinting approaches struggle to reliably detect reuse relations when the two models have different architectures or address different classification tasks (heterogeneous cases). These limitations hinder the practical use of fingerprinting methods in diverse scenarios.
@@
To address the aforementioned limitations, we propose a fingerprinting method called Neuron Functionality Analysis-based Reuse Detector (NFARD), which only requires normal test samples to detect reuse relations. This method measures the differences between models based on a simple yet effective model characterization, i.e., neuron functionality (NF). Neuron functionality has been well practiced to characterize a DNN in neural network representation studies [26, 27]. In this work, we leverage neuron functionality for copyright protection. Specifically, a neuron’s functionality is the mapping from the input space to the one-dimensional output space of the neuron, i.e., $\mathcal{X}\rightarrow\mathbb{R}$ with $\mathcal{X}$ representing the input space. [^1] Neuron functionality is the direct mapping function which is fundamentally different from neuron output statistics used in previous approaches. Recent studies [26, 27, 28] have shown that independently trained neural networks have divergent neuron functionalities. On the other hand, neural networks obtained by reuse techniques have similar neuron functionalities to those of the original model. Therefore, we can detect model reuse relations by measuring the similarity of neuron functionalities, which can be approximated only using normal samples, eliminating the need for adversarial example generation and its associated complexities. A side benefit is that the accessibility of gradient information is not required, making NFARD applicable to more reuse techniques under both white-box and black-box settings. We design specific distance metrics to measure the similarity of neuron functionality under these scenarios. Furthermore, to tackle the challenge of heterogeneous reuse detection, NFARD incorporates a linear transformation technique to extend its application scope. By constructing an optimal projection matrix via least squares optimization, we map potentially high-dimensional NF features from models with different architectures or tasks onto a common, comparable low-dimensional space. This significantly extends NFARD’s applicability to diverse real-world reuse scenarios where existing methods fall short.
@@
To enable a comprehensive evaluation of NFARD, we construct a model reuse detection benchmark named Reuse Zoo, which consists of 250 neural network models, covering seven commonly used model reuse techniques, three datasets, and four popular model architectures. We use this benchmark to evaluate NFARD and compare it with three baseline fingerprinting approaches: IPGuard [20], and two state-of-the-art methods, i.e., ModelDiff [24] and DeepJudge [23]. Extensive experimental results show that NFARD can effectively detect reuse relations in both black-box and white-box settings, achieving an F1-score of 0.984 and 1.0, respectively, and is $2\sim 99$ times faster than previous methods in terms of test suite generation.
Our main contributions are:
@@
    NFARD: A novel, adversarial example-free DNN copyright protection method based on neuron functionality analysis, applicable in both white-box and black-box settings using normal test samples.
@@
    Heterogeneous reuse detection: A linear transformation technique enabling NFARD to effectively handle reuse scenarios involving models with different architectures or classification tasks.
@@
    Reuse Zoo benchmark: A publicly available benchmark with 250 models designed for evaluating DNN reuse detection methods, covering seven common reuse techniques.
@@
    Comprehensive evaluation: Demonstrating NFARD’s superior performance, broader application scope, and efficiency compared to the state-of-the-art approaches.
@@ II-A Deep Neural Network and Neuron Functionality
We focus on classification tasks in this work. A deep neural network classifier is a function mapping from the input space to a probability distribution, i.e., $f:\mathcal{X}\to[0,1]^{m}$ where $\mathcal{X}$ is the input space and $m$ is the number of classes. The neural network includes $L$ layers whose output dimensions are $m_{1},m_{2},\cdots,m_{L}$, respectively (with $m_{L}=m$). We assume that the last layer is a fully connected layer with $m$ neurons, and its output is called the logits of the model. The final probability distribution is obtained by applying the softmax function to the logits. Here, $f$ defines the architecture (e.g., layer composition). A trained model is an instance of certain architecture $f$ with a set of parameters $\theta$, denoted as $f(\,\cdot\,;\theta)$.
In a neural network $f$, the functionality of a neuron is the mapping from the input space to the one-dimensional output space of the neuron. For the $i$-th neuron of the $k$-th layer, its functionality is defined as $f_{i}^{(k)}:\mathcal{X}\to\mathbb{R}$. Mathematically, $f_{i}^{(k)}$ is a highly non-convex function. In practice, it can be approximated by a set of normal samples $\{\bm{x}_{j}\}_{j=1}^{n}$ and characterized as a vector denoted as $\bm{\mathrm{nv}}(k,i)$:
@@
Recent work [26, 27] on neural network representation proposed a neuron functionality analysis approach to study DNN learning dynamics and interpretability. They have shown that independently trained DNN models with different weight initialization have different corresponding neuron functionalities, which is in accordance with [28]’s conclusions. That is to say, even if two DNN models $f(\,\cdot\,;\theta_{1})$ and $f(\,\cdot\,;\theta_{2})$ have the same structure, training dataset, and training setting, their corresponding neuron functionalities $\bm{\mathrm{nv}}_{1}(k,i)$ and $\bm{\mathrm{nv}}_{2}(k,i)$ can vary greatly if they have different weight initialization.
@@ II-B Model Reuse Detection for Copyright Protection
Given a victim model $f_{v}(\,\cdot\,;\theta_{v})$ and a suspect model $f_{s}(\,\cdot\,;\theta_{s})$, the task of the reuse detector is to ascertain whether $f_{s}(\,\cdot\,;\theta_{s})$ is a positive suspect model, i.e., a surrogate model derived from $f_{v}(\,\cdot\,;\theta_{v})$ via reuse methods, or a negative suspect model, independently trained or unrelated to $f_{v}(\,\cdot\,;\theta_{v})$. To facilitate this determination, the reuse detector typically employs a set of reference models $\{f_{r_{i}}(\,\cdot\,;\theta_{r_{i}})\}_{i=1}^{R}$ [21, 24, 23]. Each reference model shares the same structure as the victim model but is trained independently [^3] In practice, hyperparameter tuning process (e.g., grid search) produces multiple independently trained models that can be used as reference models., providing a reference for the detector to make the decision. This independence ensures that the reference models represent typical variations in model behavior when trained without knowledge transfer from $f_{v}(\,\cdot\,;\theta_{v})$. For brevity, we will omit the parameters $\theta_{v}$ (and $\theta_{s}$, $\theta_{r_{i}}$) when they are clear from the context or not important, and use $f_{v}$ (and $f_{s}$, $f_{r_{i}}$) to represent the model.
Depending on the access of the reuse detector to the models, the detection methods can be divided into two categories: white-box and black-box methods. In the white-box setting, the detector has access to the internals of the neural networks and the final probability output. While in the black-box setting, the detector only has access to the final probability output. In either case, we assume that a subset of the training data of the victim model is available for the reuse detector (e.g., the model owners themselves are searching for surrogate models).
Another important aspect for detection is the reuse type of the suspect model. Depending on whether the suspect model has the same structure and the same classification task as the victim model, its reuse type can be classified as homogeneous or heterogeneous. A homogeneous suspect model has the same structure as the victim model and is used for the same classification task. On the other hand, a heterogeneous suspect model has a different structure or is used for a different classification task. Detecting heterogeneous surrogate models is usually more challenging. To the best of our knowledge, in previous work, only ModelDiff [24] could handle transfer learning reuse cases, but it is not good at dealing with the other two heterogeneous reuse types. We propose a benchmark Reuse Zoo covering different types of surrogate models for DNN reuse detector evaluation, and we elaborate on the construction details in Section IV.
@@ II-C Model Reuse Methods
We consider seven commonly used DNN reuse methods, including fine-tuning, retraining, pruning, quantization, knowledge distillation, transfer learning, and model extraction.
Fine-tuning and retraining techniques train the victim model for a few epochs with a relatively small learning rate, and some weight parameters may be frozen [30] to enable a finer control. The difference between them is that retraining also re-initializes the weights of the last layer. Pruning and quantization methods are compression techniques that aim to reduce DNN size by removing trivial weights or quantizing parameters to low-bit values while minimizing the drop in model performance [31, 32]. Surrogate models obtained by these four methods have the same model structure and classification task as the victim model, so they are homogeneous models. On the other hand, knowledge distillation and model extraction train surrogate models by exploiting the victim model’s outputs [33, 9, 10], and they usually result in a different architecture; transfer learning replaces the last layer with a new one [5], and the classification task becomes different. Therefore, these three reuse techniques produce heterogeneous models with repsect to the victim model.
@@ III-A Insight into Reused DNN Models
Given a set of test inputs $X=\{\bm{x}_{1},\cdots,\bm{x}_{n}\}$, we introduce two metrics to measure the similarity of neuron functionality. The first metric is $\mathrm{ED}^{(k)}$, which is the average Euclidean distance between the $k$-th layer’s neuron vectors of two models:
$\mathrm{ED}^{(k)}(f_{v},f_{s})=\dfrac{1}{m_{k}}\sum\limits_{i=1}^{m_{k}}\lVert\bm{\mathrm{nv}}_{v}(k,i)-\bm{\mathrm{nv}}_{s}(k,i)\rVert_{2},$ (2)
where $\bm{\mathrm{nv}}_{v}(k,i)$ is the $i$-th neuron vector of the $k$-th layer of $f_{v}$, and $\bm{\mathrm{nv}}_{s}(k,i)$ is that of $f_{s}$. The second metric is the average cosine distance of the neuron vectors, denoted as $\mathrm{CD}^{(k)}$:
$\mathrm{CD}^{(k)}(f_{v},f_{s})=\dfrac{1}{m}\sum\limits_{i=1}^{m}(1-\dfrac{\bm{\mathrm{nv}}_{v}(k,i)^{\intercal}\bm{\mathrm{nv}}_{s}(k,i)}{\lVert\bm{\mathrm{nv}}_{v}(k,i)\rVert_{2}\lVert\bm{\mathrm{nv}}_{s}(k,i)\rVert_{2}}).$ (3)
Compared with $\mathrm{ED}^{(k)}$, $\mathrm{CD}^{(k)}$ focuses more on the difference in direction rather than the scale. The smaller the metrics, the more similar the two models are in terms of neuron functionality. We do not include metrics used in previous DNN representation studies (such as SVCCA similarity [26]) because they target analyzing training dynamics on hidden layers where correlation analysis is essential, and our goal is to detect reuse relations using the output layer neuron vectors.
@@
Next, we examine four victim models with different structures and learning tasks. For each victim model, we prepare 8 surrogate models by applying homogeneous reuse methods and 40 reference models that are independently trained on the same task (details of these models are available in Section IV). We compute the distance metrics $\mathrm{ED}^{(k)}$ and $\mathrm{CD}^{(k)}$ on the last layer (before the softmax function) using 1000 normal samples, and the results are presented in Figure 1, where surrogate models and reference models are plotted in orange and blue, respectively. We can see that reference models have larger distance values and are concentrated in the upper right corner of the plane, while surrogate models have smaller distance values and are scattered on the lower left part. Such differences indicate that homogeneous surrogate models are more similar to victim models than independently trained models regarding neuron functionality, which is exactly the principle of our proposed method.